### v1

# Intentional Gesture: Deliver Your Intentions with Gestures for Speech

Abstract
When humans speak, gestures help convey communicative intentions, such as adding emphasis or describing concepts. However, current co-speech gesture generation methods rely solely on superficial linguistic cues (e.g. speech audio or text transcripts), neglecting to understand and leverage the communicative intention that underpins human gestures. This results in outputs that are rhythmically synchronized with speech but are semantically shallow. To address this gap, we introduce Intentional-Gesture, a novel framework that casts gesture generation as an intention-reasoning task grounded in high-level communicative functions. First, we curate the InG dataset by augmenting BEAT-2 with gesture-intention annotations (i.e., text sentences summarizing intentions), which are automatically annotated using large vision-language models. Next, we introduce the Intentional Gesture Motion Tokenizer to leverage these intention annotations. It injects high-level communicative functions (e.g., intentions) into tokenized motion representations to enable intention-aware gesture synthesis that are both temporally aligned and semantically meaningful, achieving new state-of-the-art performance on the BEAT-2 benchmark. Our framework offers a modular foundation for expressive gesture generation in digital humans and embodied AI. Project Page: https://andypinxinliu.github.io/Intentional-Gesture
Figure 1: We present Intentional Gesture, a novel framework for intention-controllable gesture generation. Our method models latent communicative functions from speech and grounds motion generation in these inferred intentions.
1 Introduction
Gestures accompany speech to convey emotions, emphasize ideas, and facilitate comprehension [6, 30, 3], forming an essential part of human communication. As AI advances, enabling virtual avatars to produce expressive gestures will be critical for digital avatar construction [32, 29, 9, 27, 28].
Recent works [35, 16, 13, 38] have explored full-body gesture generation conditioned on speech audio or text transcripts. However, these approaches largely treat raw linguistic input as sufficient semantics, overlooking the deeper communicative intentions—such as emphasis, deixis, and affirmation—that implicitly govern gesture behavior. Recovering these latent functions is essential for generating semantically coherent non-verbal expressions.
Early systems [19, 26, 20] attempted to derive these deeper communicative functions through rule-based templates, but such approaches lack scalability and domain adaptability. Inspired by the linguistic–gesture link in psycholinguistics, we reframe gesture modeling as a reasoning task: the model first understands the reasons or communicative intentions behind the speech, and then treats gestures as their downstream realizations.
To achieve this, we first augment the BEAT-2 dataset [14] with structured annotations of inferred communicative functions via VLMs (Vision-Language-Models), creating a large-scale InG (Intention-Grounded) dataset. This enriched corpus links speech, inferred intentions, and corresponding gesture realizations, enabling intention-aware gesture generation for the first time.
Leveraging the obtained intention annotations, we propose a CLIP-like gesture understanding model. It learns joint representations of speech rhythm, inferred intentions, and motion dynamics via a hierarchical contrastive alignment framework, enabling it to discern both low-level rhythmic synchronization and high-level communicative goals.
Building upon this understanding model, we introduce the Intentional Gesture Tokenizer, a novel quantization model that directly embeds intention semantics into the motion representation space. Unlike prior works that discretize body parts independently without semantic grounding [21, 15], our tokenizer processes global body motion holistically and supervises latent representations using intention-aware motion features. This design ensures that discrete motion tokens encode not only fine-grained motion patterns but also communicative meaning. With the proposed tokenizer, our approach generates gestures that are not only temporally synchronized but also semantically expressive and interpretable, advancing human-avatar communication. Our contributions can be summarized as follows:
  • We formulate gesture generation as an intention-grounded reasoning task, create the InG (Intention-Grounded) dataset by leveraging VLMs to infer communicative functions from speech and augmenting BEAT-2 with structured annotations.
  • We introduce the Intentional Gesture Tokenizer, which discretizes global body motion while embedding intention semantics into the latent space through semantic supervision.
  • We demonstrate that our method produces gestures that are not only temporally aligned with speech but also semantically meaningful, achieving improved realism and interpretability in human-computer interaction.
2 Related Works
Co-speech Gesture Generation. Existing works on co-speech gesture generation typically use skeleton- or joint-level pose representations. Several methods [17, 7, 36, 12, 42, 16] learn hierarchical semantics or apply contrastive learning to align audio and gesture embeddings. HA2G [17] constructs multi-level audio-motion embeddings, while TalkShow [38], CaMN [14], and EMAGE [13] introduce large-scale datasets for joint face-body modeling with GPT-style decoding. More recent models such as MambaTalk [35], DiffSHEG [5], and GestureLSM [15] focus on efficient flow matching and spatiotemporal modeling for quality-speed tradeoffs. Semantic Gesticulator [43] and RAG-Gesture [21] leverage LLMs to retrieve discourse-relevant gestures as references. In contrast, we explore the modeling of communicative intentions as explicit semantics to control gesture generation.
Vision Tokenization for Generation. In visual generation, tokenization encodes raw pixels into compact representations [34, 25]. Vector-quantization [34] enable discrete latent spaces compatible with autoregressive generation [33, 31, 4, 40]. Recent studies like REPA [41] and Re.vs.Gen [37] show that aligning generation representation with understanding improves synthesis quality. We further introduce the Intentional Gesture Tokenizer, supervising motion tokenization with intention-aligned semantics, and demonstrate this strategy significantly enhances gesture generation quality.
Figure 2: Overview of the annotation pipeline. Sentence-level segments and word-aligned keyframes are provided to the VLM, which generates (1) motion descriptions across body regions, (2) communicative function labels, (3) gesture behavior mappings, and (4) inferred intentions. This structured annotation enables scalable creation of a semantically grounded, visually aligned gesture dataset.
3 Intentional Gesture Dataset (InG)
Gestures systematically reflect semantic, pragmatic, and rhetorical structures [19, 20, 26]; for example, affirmation to nodding and comparison to lateral gestures aligned with spatial metaphors. Modeling these communicative functions is critical for generating expressive gestures. Marsella et al. [19] introduced a rule-based system that derives communicative functions from syntactic and lexical patterns, mapping them to gesture classes via hand-crafted rules. While effective, such systems rely on fixed rules and limited dictionaries, restricting scalability and domain coverage. Building on these insights, we propose a scalable pipeline using VLMs to automatically infer communicative functions and map them to gestures, enabling large-scale, data-driven semantic gesture modeling.
3.1 Dataset Construction
We introduce InG (Intention-Grounded Gestures), a dataset that augments BEAT-2 [14] with structured, intention-aware annotations linking linguistic functions to gesture behaviors. Unlike prior datasets focused on motion fidelity or prosody alignment, InG enables intention-controllable gesture generation. We employ a two-stage VLM-based strategy, illustrated in Fig. 2.
Annotation Protocol. For training data, the VLM receives sentence context and word-aligned keyframes, allowing inference of communicative functions, motion descriptions, and speaker intent. For testing data, annotations are generated from transcripts only to prevent leakage from observed gestures. Detailed prompting strategies and validations are provided in the Appendix.
Sentence-Level Segmentation and Keyframe Extraction. To prepare BEAT-2 for intention annotation, we render each sequence into frame-wise visual data. Speech sequences are segmented into sentence-level units using Whisper-X [1], which restores punctuation and provides word-level alignment. Boundaries are identified via punctuation, and timestamps are extracted. For each word, we uniformly sample one keyframe within its temporal span to capture the pose at word-level granularity.
Prompt-Based Gesture Intention Annotation. Given the segmented utterances and word-level keyframes, we design a structured prompt to guide GPT-4o-mini in multi-stage annotation: (1) Motion Analysis: describing body poses across regions (head, arms, fingers, torso); (2) Function Derivation: identifying communicative functions (e.g., Emphasis, Negation, Deixis); (3) Gesture Behavior Mapping: associating functions with gesture forms; (4) Inferred Intention: synthesizing how verbal and nonverbal cues reflect speaker goals. This structured approach enables scalable creation of semantically and physically grounded gesture annotations.
Figure 3: (a) Distribution of communicative function types in InG dataset. (b-d) Correlations between function count and utterance-level features: word count, duration, and speech rate. Function count correlates positively with utterance length and duration, but minimally with speech rate.
3.2 Dataset Statistics and Analysis
Our InG dataset includes 34,641, 3,598, and 9,674 annotated utterances for train/val/test splits, respectively, each enriched with motion-grounded or intent-inferred communicative functions. We defer the specific statistics of the distribution of 16 function types, covering pragmatic categories such as Emphasis, Deixis, Mental State, and Process in the Appendix.
As shown in Fig. 3 (a), Emphasis (21.7%) and Deixis (20.1%) are the most prevalent functions, followed by Mental State (14.4%) and Process (11.0%). We further analyze correlations between linguistic features and function density (Fig. 3, b-d). Function counts correlate positively with utterance length (r = 0.18–0.20) and duration (r = 0.16–0.21), but show minimal correlation with speech rate (r $\approx$ 0.00–0.01). This suggests that communicative function density scales with the informational content rather than speaking speed, aligning with psycholinguistic findings.
While function derivations follow a well-defined ontology, gesture behavior mappings and inferred intentions exhibit open-ended variability across speakers and contexts. To ensure interpretability, our prompt design grounds mappings in established literature (McNeill [20], Kendon [26]). Additional statistics, qualitative examples and function co-occurrence analysis are provided in Appendix.
Figure 4: Pairwise preference for VLM vs. human annotations across three evaluation settings. Blue: train-style prompting (with motion); Orange: test-style (transcript only). Red dashed line indicates chance (50%).
Annotation Validation Study. We evaluate annotation quality through a pairwise human preference study on 100 randomly sampled utterances. Each utterance is annotated using: (1) our train-style VLM protocol (motion input), (2) our test-style VLM protocol (transcript only), and (3) a human-written baseline. For each utterance, two comparisons are created: VLM (train-style) vs. human, and VLM (test-style) vs. human.
Three raters independently judge each comparison across three criteria: E1: Intent Alignment (which better captures communicative intent), E2: Gesture Relevance (which provides more plausible gestures), and E3: Overall Preference (clarity and function-gesture alignment).
Final preferences are determined by majority vote. Shown in Fig. 4, both VLM protocols outperform human annotations, especially on Intent Alignment and Overall Preference. Inter-rater agreement averaged 0.76 Fleiss’ $\kappa$, indicating substantial consistency. These results validate the effectiveness of our structured prompting and the scalability of VLM-based annotation. Additional validation details, including hallucination audits and mapping coherence analysis, are in Appendix A.3–A.6.
Figure 5: Overview of our intentional gesture generation framework. Left: H-AuMoCLIP learns a hierarchical joint embedding of motion, audio, and intention. Transcript embeddings (BERT) aligned via CTC serve as queries in a cross-attention module with intention embeddings as keys/values. The resulting semantic features are concatenated with wav2vec2 audio features for contrastive learning. Right: Motion is quantized via a multi-codebook VQ module and supervised by semantic features from H-AuMoCLIP, enabling expressive and controllable gesture generation.
4 Intentional Gesture Generation
4.1 Intention Understanding
To leverage textual intention annotations for co-speech gesture synthesis (akin to text-to-image generation), we propose a CLIP-like understanding model, H-AuMoCLIP, for gestures to better encode textual intention features. We build this Intention-Audio-Motion CLIP on top of TANGO’s [12] audio-motion alignment model (AuMoCLIP) by explicitly incorporating communicative intent.
Intentional Semantic Fusion. AuMoCLIP trains its joint embedding space using a two-tower contrastive learning framework, where its audio tower typically combines acoustic features (e.g., wav2vec2) with BERT-encoded speech transcripts. Building upon this, we introduce a dedicated pathway, Intentional Semantic Fusion, for merging intention semantics to create a richer, intention-aware joint embedding space. Our Intentional Semantic Fusion mechanism fuse transcript and intention embeddings via a linguistically grounded mechanism. Following TANGO, we align transcript tokens to the audio timeline using a CTC-based model, and encode both the transcript and annotated intent with separate BERT encoders. In addition to TANGO, the aligned transcript embeddings serve as queries in a cross-attention module, with intention embeddings as keys and values. The output captures context-specific communicative cues, which are then concatenated with wav2vec2 audio features to form the final high-level representation. This composite embedding is used in contrastive training to jointly align audio, motion, and intent. Note that we retain TANGO’s low-level contrastive training setup to align audio-motion features rhythmic level as well.
Functionality. As in Fig.5, the resulting encoders improve generation by serving as two roles: (1) the motion encoder provides semantic supervision for gesture tokenization (Sec. 4.2), and (2) the audio encoder conditions gesture generation with both rhythmic and intentional signals (Sec. 4.3).
4.2 Intentional Gesture Tokenizer
Prior works [21, 15, 13, 35] often model different body regions using separate quantizers, maintaining distinct encoders, decoders, and codebooks per body part. While this tokenizer encourages local disentanglement, it introduces three limitations: (1) high training and inference cost, (2) poor global body coherence, and (3) weak connection to speech-level semantics. To address these issues, we propose an Intentional Gesture Tokenizer, which introduces two core innovations: (1) learning disentangled latent factors within a unified global body representation through multi-codebook quantization, and (2) embedding explicit intention semantics into the motion representation through semantic supervision.
Architecture Design. While CNNs are effective at capturing fine-grained local details [23, 2, 15, 13], Transformers offer greater scalability with reduced inductive bias [8]. To leverage the strengths of both, we adopt a hybrid VQ-tokenizer architecture: the encoder comprises CNN blocks that downsample the input gesture motion by a factor of 4, followed by 1 Transformer layers and a vector quantizer that outputs discrete latent codes. The decoder mirrors this structure in reverse: Transformer layers followed by CNN blocks that upsample the features to reconstruct the gesture motion.
Latent Multi-Codebook Quantization. Instead of separate body-specific quantizers, we discretize the global motion latent $f\in\mathbb{R}^{d}$ using a set of $n$ independent codebooks. The latent is partitioned into $n$ chunks $\{f_{1},f_{2},\dots,f_{n}\}$, each quantized separately:
$\hat{f}=\text{Concat}\left(\mathcal{Q}(Z_{1},f_{1}),\dots,\mathcal{Q}(Z_{n},f_{n})\right),$ (1)
where $\mathcal{Q}$ denotes vector quantization. This design maintains global body context while allowing structured disentanglement to emerge across codebooks during training. Unlike prior methods, our tokenizer processes the full body motion jointly, enabling better modeling of coordination patterns necessary for communicative gestures.
Intentional Semantic Supervision. To embed contextual intent into motion representations, we supervise the quantized latents using features from the pretrained H-AuMoCLIP motion encoder (Sec. 4.1). A linear projection is applied to the quantized output $Z$ to match the dimension of reference features $F$, and we compute a temporal cosine margin loss:
$\mathcal{L}_{\text{sem}}=\frac{1}{T}\sum_{t=1}^{T}\text{ReLU}\left(1-\frac{z^{\prime}_{t}\cdot f_{t}}{\|z^{\prime}_{t}\|\|f_{t}\|}\right),$ (2)
where $z^{\prime}_{t}$ and $f_{t}$ are projected quantized latents and motion encoder features at timestep $t$. This supervision enforces semantic alignment with high-level intention cues derived from speech.
Training Objective. Our full training objective includes: (i) a reconstruction loss $\mathcal{L}_{\text{R}}$, (ii) a vector quantization loss $\mathcal{L}_{\text{VQ}}$, and (iii) the proposed semantic supervision loss $\mathcal{L}_{\text{sem}}$:
$\mathcal{L}=\mathcal{L}_{\text{R}}+\lambda_{\text{VQ}}\mathcal{L}_{\text{VQ}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}}.$ (3)
We empirically set $\lambda_{\text{sem}}=1$ and $\lambda_{\text{VQ}}=0.25$. This balanced formulation ensures that quantized motion tokens preserve both fine-grained fidelity and semantic coherence across time.
4.3 Gesture Generation
We then leverage Sec.4.1 and Sec.4.2 to enhance gesture generation. We adapt a state-of-the-art gesture generation model, GestureLSM [15], by replacing its original audio encoder and motion tokenizer. Specifically, we replace GestureLSM’s audio encoder with our intent-aware audio encoder, which extracts hierarchical representations capturing both rhythmic audio features and communicative intentions. Furthermore, we replace its RVQ-based tokenizer with our Intentional Gesture Tokenizer. Additional architectural details of the original GestureLSM are deferred to the Appendix.
5 Experiments
We conduct all experiments on BEAT2 [13], comprises 60 hours of high-quality SMPL-based gesture data collected from 25 speakers (12 female, 13 male). The dataset contains 1,762 sequences, each with an average duration of 65.66 seconds, where each sequence captures responses to daily inquiries. For consistency, we adopt the train-validation-test split protocol defined in EMAGE [13].
5.1 Quantitative Comparisons
Metrics. We evaluate Fréchet Gesture Distance (FGD)[39] for distributional similarity between ground truth and synthesized gestures. Diversity (Div.)[10] is measured by the average L1 distance across gesture clips, while speech-motion synchronization is assessed via Beat Constancy (BC) [11].
Evaluation Results. We evaluate both quality and generalization in three settings: single-speaker, all-speaker, and zero-shot (training on 20 speakers and testing on 5 unseen speakers). As shown in Table 1, our method significantly outperforms baselines across all settings. These gains are attributed to our intentional motion representation and intention control, which reduce unnatural gesture patterns. More detailed comparisons are provided in the Appendix. To validate runtime efficiency, we generate intention descriptions from transcripts via an API prior to model execution. Our method retains the efficiency of GestureLSM [15], while improving performance through a unified motion representation, and achieves real-time generation at 30.4 FPS on an H100 GPU.
Figure 6: The subjective Comparisons. Compared with other methods, Intentional-Gesture presents more natural and coherent motion patterns to represent specific words or phrases (highlighted in red).
Table 1: Comparison with state-of-the-art methods trained on BEAT-2. We demonstrate superior performance, especially when generalizing across multiple speaker identities.
1 Speaker All Speakers Novel 5 speakers
[…]
GT $0.703$ $11.97$ $0.477$ $7.29$ $0.671$ $8.93$
CaMN [14] $0.604$ $0.676$ $9.97$ $0.512$ $0.200$ $5.58$ $0.812$ $0.563$ $6.71$
EMAGE [13] $0.570$ $0.793$ $11.41$ $0.692$ $0.284$ $6.06$ $0.936$ $0.643$ $7.47$
Audio2Photoreal [22] $1.02$ $0.550$ 12.47 $0.849$ $0.326$ $6.24$ $1.012$ $0.464$ $5.97$
RAG-Gesture [21] $0.879$ $0.730$ $12.62$ $0.447$ 0.471 $9.03$
GestureLSM [15] $0.408$ $0.714$ $13.24$ $0.446$ $0.525$ $9.23$ $0.664$ $0.621$ $10.45$
Ours 0.379 0.690 $11.00$ 0.256 $0.534$ 6.68 0.441 0.686 9.39
5.2 Qualitative Comparisons
Evaluation Results As depicted in Figure 6, our approach generates gestures that exhibit improved rhythmic alignment and a more natural appearance compared to existing methods. For example, when conveying the phrase “helpful for me”, our method directs the subject to extend left hand forward, effectively representing the intention of an explanation. In contrast, competing methods fail to capture this nuance, often generating static or unnatural poses where one or both arms remain down.
Table 2: Subjective evaluation shown as Mean Opinion Scores (MOS).
Methods MOS1 MOS2 MOS3
EMAGE 2.01 2.42 2.31
CAMN 1.34 2.23 2.14
GestureLSM 3.43 3.61 3.48
Ours 3.76 4.11 3.92
User Study. We conducted a user study with 20 participants and 160 video samples, 40 from each of GestureLSM [15], EMAGE [13], CAMN [14], and Ours. Each participant viewed the videos in a randomized order and rated them on a scale of 1 (lowest) to 5 (highest) based on three criteria: (1) realness, (2) speech-gesture synchrony, and (3) smoothness. As shown in Tab.2, Our method outperforms other methods across all criteria, achieving higher Mean Opinion Scores (MOS).
6 Ablation Studies and Analysis
We explore the hierarchical audio-intention and motion alignment, the design of tokenizer and generator in this section. We defer the further analysis of the quality of annotation effects on learning, model generalization ability, sequence length effect, and quantizer analysis to Appendix E.
Figure 7: Hierarchical Alignment Analysis. Retrieval-based evaluation of H-AuMoCLIP across low-level (rhythmic) and high-level (semantic) features. We report Recall@1/10 under per-batch and global settings for both known and unknown speakers. Incorporating intentional semantics improves retrieval performance, especially for high-level alignment and with few training speakers. Increasing speaker diversity further enhances generalization in high-level retrieval.
6.1 Evaluation of Hierarchical AuMoCLIP
We evaluate H-AuMoCLIP using retrieval-based metrics that measure how well audio features align with corresponding motion features at both low and high levels. We report Recall@1 and 10 under two settings: per-batch (within batch of 128) and global (across the entire test set).
Low-level retrieval evaluates frame-wise alignment by computing cosine similarity between audio and motion features at each timestep, within a local temporal window $(i\pm 16)$. High-level retrieval evaluates sequence-level alignment using mean-pooled audio and motion features for full-clip matching. Please see more details in TANGO [12]. We evaluate generalization to both seen (in-domain) and unseen (out-of-domain) speakers, assessing robustness across speaker identities.
Intention-Aware Alignment Improves Performance. Incorporating intentional semantics consistently improves retrieval, especially under low-data regimes (1–5 speakers) (Fig. 7). High-level alignment shows larger gains compared to low-level retrieval, confirming that explicit intention modeling better captures sequence semantics.
Speaker Diversity Enhances Generalization. Training on multiple speakers significantly improves generalization to unseen speakers. Even with the same total training data, models trained on more speaker identities achieve better high-level retrieval, suggesting that speaker diversity encourages learning more robust gesture-speech mappings, as shown in Appendix E.
Low-Level Alignment Remains Stable. In contrast, low-level beat synchronization metrics show minimal gains from increased data or speaker diversity, indicating that rhythm alignment depends more on prosodic regularities than on inter-speaker variation.
Figure 8: Tokenizer ablation. We perform both global and local grid searches to study the effects of codebook design choices on rFGD (×$\times 10^{3}$). (a)–(c): Global sweeps varying one factor at a time; (d): Local grid search over code count and dimension. All results confirm consistent trends: 8 codebooks, a code dimension of 32, and 8192 codes yield optimal or near-optimal performance.
Table 3: Quantitative Ablation Results. (Left) Quantizer ablation comparing design variants and levels of intentional supervision. (Right) Gesture generation improvements with architectural modules, semantic controls and sensitivity analysis of the whole generator model.
[…]
EMAGE (VQ) 0.0867 34.35 0.3643 0.512 0.623 7.76
RAG-Gesture (Contin.) 0.0423 - 0.448 0.423 0.625 8.64
ProbTalk (PQ) 0.0283 53.24 0.3283 0.451 0.656 7.96
GestureLSM (RVQ) 0.0012 27.56 0.1557 0.314 0.527 8.64
Ours (Multi-VQ)
only transformer 0.0021 97.25 0.3043 0.332 0.474 8.01
only CNN 0.0013 98.43 0.2656 0.348 0.671 7.96
Optimal Quantizer Design 0.0011 98.84 0.2468 0.315 0.512 7.67
+Low-level Supervision 0.0014 96.72 0.2566 0.295 0.477 5.96
+High-level (seq) 0.0014 96.68 0.2512 0.286 0.614 6.21
+High-level (temporal) 0.0012 97.17 0.2488 0.256 0.534 6.68
[…]
GestureLSM [15] 0.446 0.525 9.23
+ DiT [24] 0.354 0.523 6.63
+ Intention Condition 0.339 0.545 6.44
+ H-AuMoCLIP 0.314 0.527 8.64
+ Intentional-Tokenizer 0.256 0.534 6.68
Sensitivity Analysis
Re-summarized intent 0.256 0.533 6.66
Unmatched intent 0.284 0.521 6.47
Window = 0.7 0.257 0.533 6.64
Window = 0.5 0.256 0.533 6.67
w/o intent text 0.281 0.523 6.54
6.2 Evaluation of Intentional Gesture Tokenizer
We evaluate our intentional gesture tokenizer using reconstruct FGD (rFGD), Codebook Utility, and L1 error, with generation metrics reported in Sec. 5.1. We defer further analysis in Appendix.
Tokenizer Comparison. We leverage the same generator design in our framework with different latent representation from various tokenizer designs for this comparison. As in Tab. 3 Left, our Intentional Gesture Tokenizer achieves a significantly lower rFGD and higher diversity compared to previous tokenizers [13, 18, 21, 15].
Design Analysis. We ablate design choices of the tokenizer, including the number of codebooks, code dimension, and code size. Fig. 8 shows that (1) 8 codebooks outperform fewer or more, balancing representational capacity and model compactness; (2) a code dimension of 32 achieves the best trade-off between expressiveness and compression; and (3) increasing code size improves rFGD up to 8192 codes, with diminishing returns beyond. These trends are consistent across global and local grid searches. For architecture design, we discover CNN presents better reconstruction quality, but the transformer presents better generation FGD. Our hybrid design takes the advantage of both variants.
Effect of Intentional Semantic Supervision. We design three supervision strategies using H-AuMoCLIP features: (a) CNN-layer supervision for low-level spatial detail, (b) transformer-layer with mean pooling for high-level sequence supervision, and (c) transformer-layer with frame-level temporal supervision. As shown in Table 3 (Left), integrating intentional supervision into motion token learning significantly improves generation quality—both in FGD and diversity—while maintaining strong reconstruction fidelity (L1, rFGD). Compared to CNN-only supervision, transformer-based supervision better captures communicative semantics, validating the benefit of higher-level grounding. Notably, frame-level temporal supervision achieves the best overall FGD and diversity, demonstrating that temporal alignment of intention and motion is key to learning expressive gesture representations.
6.3 Evaluation of Gesture Motion Generation
Design Analysis. Replacing spatial-temporal attentions in GestureLSM with self-cross attention and DiT [24] based architecture improves the generation performance on all-speaker setting. We further investigate how intention representations impact gesture generation. As a baseline, we condition the generator on sentence-level BERT embeddings, which provides only marginal improvement (Table 3, Right). Replacing these with intention embeddings derived from H-AuMoCLIP—aligned to audio and semantic context—yields substantial gains in FGD and diversity.
Sensitivity Analysis. We evaluate the robustness of our model under noisy or mismatched intention inputs. Using LLM-rewritten paraphrases of intention descriptions yields negligible differences in FGD and diversity, suggesting that the model is not sensitive to surface-level textual variation. When using temporally unmatched or missing intention text, FGD degrades slightly (by 0.025–0.03), and synchronization weakens marginally, but the model remains stable overall. We attribute this to our alignment mechanism: we extract intention-aligned segments based on transcript word timestamps and apply a temporal overlap threshold (0.5 or 0.7) to guide supervision.
7 Conclusion
We present Intentional Gesture, a novel framework that formulates gesture generation as an intention reasoning task grounded in high-level communicative functions. We first curate the InG dataset by augmenting BEAT-2 with structured annotations of latent intentions—such as emphasis, comparison, and affirmation—extracted via large language models. Then, we introduce the Intentional Gesture Tokenizer, which discretizes global motion while embedding intention semantics into the representation space through targeted supervision. By bridging speech semantics and motion behaviors, our method produces gestures that are both temporally synchronized and semantically meaningful. Experiments on BEAT-2 demonstrate strong improvements in gesture quality, offering a controllable and modular foundation for expressive gesture generation in digital humans and embodied AI.